OpenStudy (anonymous):

@shamil98 What is the value of the expression ? x = 3 and y = 3. A. –3 B. 1 C. 9 D. 18

OpenStudy (shamil98):

Value? I'm not sure what this is asking..

OpenStudy (shamil98):

Ah.

OpenStudy (shamil98):

Plugging your numbers in.. \[\huge \frac{ 4(3) - 3 }{ 2(3) + 3 } = \frac{ 9 }{ 9 } = 1\]

OpenStudy (anonymous):

so its 1

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

The formula for the area of a triangle is A = ½bh where b is the base and h is the height. What is the area of the triangle shown here? right triangle A. 72 in.2 B. 90 in.2 C. 180 in.2 D. 360 in.2